### v1

@@ 5 Concluding remarks
In this work, we have investigated stochastic variance reduced gradient (SVRG) and a regularized variant (rSVRG) for solving linear inverse problems in Hilbert spaces. We have established the regularizing property of both SVRG and rSVRG. Under the source condition, we have derived convergence rates in expectation and in the uniform sense for (r)SVRG. These results indicate the optimality of SVRG for nonsmooth solutions and the built-in regularization mechanism and optimality of rSVRG. The numerical results for three linear inverse problems with varying degree of ill-posedness show the advantages of rSVRG over both standard SVRG and Landweber method. Note that both SVRG and rSVRG depend on the knowledge of the noise level. However, in practice, the noise level may be unknown, and certain heuristic techniques are required for their efficient implementation, e.g., as the a priori stopping rule or constructing the approximate operator $A$. We leave this interesting question to future works.
@@
The first inequality can be found in [13, Lemma 3.4]. To show the second inequality, let $\mathrm{Sp}(P)$ be the spectrum of $P$. Then there holds
@@
$\|B^{\frac{1}{2}}(I-P^{t})P^{k}\|\leq\|B\|^{\frac{1}{2}-\epsilon}\|B^{\epsilon}P^{\frac{k+t}{2}}\|\|(I-P^{t})P^{\frac{k-t}{2}}\|\leq 2^{1+\epsilon}\epsilon^{\epsilon}c_{0}^{-\epsilon}\|B\|^{\frac{1}{2}-\epsilon}t(k+t)^{-(1+\epsilon)}.$
@@
$\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=$ $\mathbb{E}[\|R(B-A_{i_{j}}^{*}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=\mathbb{E}[\|RA_{i_{j}}^{*}A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]-\|RB\Delta_{j}^{\delta}\|^{2}$
$\leq$ $L\|R\|^{2}\mathbb{E}[\|A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=L\|R\|^{2}\frac{1}{n}\sum_{i=1}^{n}\|A_{i}\Delta_{j}^{\delta}\|^{2}=n^{-1}L\|R\|^{2}\|A\Delta_{j}^{\delta}\|^{2}.$
@@
$\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=\mathbb{E}[\|RA^{*}(n^{-1}A-b_{i_{j}}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]\leq\mathbb{E}[\|RA^{*}\|^{2}\|(n^{-1}A-b_{i_{j}}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
$=$ $n\|RB^{\frac{1}{2}}\|^{2}\big(\mathbb{E}[\|b_{i_{j}}A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]-\|n^{-1}A\Delta_{j}^{\delta}\|^{2}\big)\leq n\|RB^{\frac{1}{2}}\|^{2}\mathbb{E}[\|A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]\leq\|RB^{\frac{1}{2}}\|^{2}\|A\Delta_{j}^{\delta}\|^{2}.$
These estimates and the inequality $\|RN_{j}\Delta_{j}^{\delta}\|^{2}\leq n\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$ complete the proof. ∎
The proof of Theorem 3.1 is lengthy and technical, and requires several technical lemmas. The first lemma provides bounds on the bias and variance components of the weighted successive error $A\Delta_{k}^{\delta}$ in terms of the iteration index.
@@
Let $k=KM+t$ with $K\geq 0$ and $1\leq t\leq M-1$. Similar to the proof of Lemma 3.2, for the bias $\|\mathbb{E}[A\Delta_{KM+t}^{\delta}]\|$, by the definitions of $\zeta$, $B$ and $\tilde{\phi}^{t-1}$, the identity
@@
Next let $S_{t,j}=c_{0}\|A(P^{t}-I)P^{KM-1-j}N_{j}\Delta_{j}^{\delta}\|$ and $T_{t,j}=c_{0}\|AP^{KM+t-1-j}N_{j}\Delta_{j}^{\delta}\|$. Then for the variance, when $K\geq 1$, by Lemma 3.1 and the identity $\mathbb{E}[\langle N_{i},N_{j}\rangle|\mathcal{F}_{i}]=0$ for any $j>i$, we have
@@
with ${\rm I}_{1}=\sum_{j=1}^{(K-1)M+t-2}\mathbb{E}[S_{t,j}^{2}],\quad{\rm I}_{2}=\sum_{j=(K-1)M+t-1}^{KM-1}\mathbb{E}[S_{t,j}^{2}]\quad\mbox{and}\quad{\rm I}_{3}=\sum_{j=KM+1}^{KM+t-1}\mathbb{E}[T_{t,j}^{2}].$
@@
$\mathbb{E}[S_{t,j}^{2}]$ $\leq c_{0}^{2}L\|B^{\frac{1}{2}}(P^{t}-I)P^{KM-1-j}\|^{2}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\leq c_{0}^{2}L\|B\|\|(P^{t}-I)P^{KM-1-j}\|^{2}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
@@
${\rm I}_{2}\leq$ $c_{0}^{2}L\|B\|\sum_{j=(K-1)M+t-1}^{KM-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\quad\mbox{and}\quad{\rm I}_{3}\leq c_{0}^{2}L\|B\|\sum_{j=KM+1}^{KM+t-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}].$
Meanwhile, by the commutativity of $B,P$ and Lemma A.1 with $\epsilon=\frac{1}{2}$, we get
@@
${\rm I}_{2}\leq$ $\frac{1}{2}c_{0}L\sum_{j=(K-1)M+t-1}^{KM-1}(KM-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
${\rm I}_{3}\leq$ $\frac{1}{2}c_{0}L\sum_{j=KM+1}^{KM+t-1}(KM+t-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}].$
@@
$\mathbb{E}[\|A\Delta_{t}^{\delta}-\mathbb{E}[A\Delta_{t}^{\delta}]\|^{2}]=$ $\sum_{j=1}^{t-1}\mathbb{E}[T_{t,j}^{2}]\leq\left\{\begin{aligned} c_{0}^{2}L\|B\|\sum_{j=1}^{t-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],\\ \frac{1}{2}c_{0}L\sum_{j=1}^{t-1}(t-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}].\end{aligned}\right$.
Then combining the preceding estimates with $\|B\|=n^{-1}\|A\|^{2}$ and $t\leq M-1$ gives the estimate (A.2). Finally, when $K\geq 1$, by Lemma 3.1 and the triangle inequality, we derive
$\|\Delta_{KM+t}^{\delta}-\mathbb{E}[A\Delta_{KM+t}^{\delta}]\|\leq$ $\sum_{j=1}^{KM-1}S_{t,j}+\sum_{j=KM+1}^{KM+t-1}T_{t,j}.$
@@
Combining these estimates with $t\leq M-1$ gives the estimate (A.3). ∎
The next lemma gives several basic estimates on the following summations
$\overline{\Phi}_{j_{1}}^{j_{2}}(i,r)$ $=\sum_{j=j_{1}}^{j_{2}}(j_{2}+i-j)^{-r}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\quad\mbox{and}\quad\Phi_{j_{1}}^{j_{2}}(i,r)=\sum_{j=j_{1}}^{j_{2}}(j_{2}+i-j)^{-r}\|A\Delta_{j}^{\delta}\|.$
@@
where $c_{K,1}=1+\frac{2}{K}$, $c_{K,2}=2+\frac{3}{K+1}+\frac{6\ln M}{(K+1)^{2}}$ and $c_{K,\epsilon}=\frac{e^{-1}+1}{\epsilon}+\frac{2^{\epsilon}\ln M}{(K+1)^{\epsilon}}$ with $K=[k/M]$, $\lim_{K\to\infty}c_{K,1}=1$, $\lim_{K\to\infty}c_{K,2}=2$, and $\lim_{K\to\infty}c_{K,\epsilon}=(e^{-1}+1)\epsilon^{-1}$.
@@
Let $k=KM+t$ with $K\geq 1$ and $t=1,\cdots,M-1$. Then there holds the inequality:
$(k-1)^{-1}\leq c_{K,1}(k+M)^{-1}.$ (A.12)
@@
$\overline{\Phi}_{k_{\rm c}+1}^{k-1}(0,0)$ $=\sum_{j=k_{\rm c}+1}^{k-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\leq c^{2}M(k-1)^{-2}\leq c_{K,1}^{2}c^{2}M(k+M)^{-2},$
$\Phi_{k_{\rm c}+1}^{k-1}(0,0)$ $=\sum_{j=k_{\rm c}+1}^{k-1}\|A\Delta_{j}^{\delta}\|\leq cM(k-1)^{-1}\leq c_{K,1}cM(k+M)^{-1}.$
@@
$\overline{\Phi}_{k_{\rm c}+1}^{k_{M}-1}(0,1)+\overline{\Phi}_{k_{M}+1}^{k-1}(0,1)$
$=$ $\sum_{j=k_{\rm c}+1}^{k_{M}-1}(k_{M}-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]+\sum_{j=k_{M}+1}^{k-1}(k-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]$
$=$ $\sum_{j=0}^{M-t}j^{-1}\mathbb{E}[\|A\Delta_{k_{M}-1-j}^{\delta}\|^{2}]+\sum_{j=0}^{t-2}j^{-1}\mathbb{E}[\|A\Delta_{k-1-j}^{\delta}\|^{2}]$
$\leq$ $c^{2}\bigg(\sum_{j=0}^{M-t}j^{-1}(k_{M}-1-j+M)^{-2}+\sum_{j=0}^{t-2}j^{-1}(k-1-j+M)^{-2}\bigg)$
$\leq$ $c^{2}(k-1)^{-2}{\rm I},\quad\mbox{with }{\rm I}=\sum_{j=0}^{M-t}j^{-1}+\sum_{j=0}^{t-2}j^{-1},$
where $\rm I$} is bounded by
${\rm I}\leq 4+\ln(M-t)+\ln t\leq 4+2\ln\tfrac{M}{2}=4-2\ln 2+2\ln M\leq 3+2\ln M.$
@@
$\overline{\Phi}_{k_{\rm c}+1}^{k_{M}-1}(0,1)+\overline{\Phi}_{k_{M}+1}^{k-1}(0,1)\leq c_{K,1}^{2}c^{2}(3+2\ln M)(k+M)^{-2}.$
@@
$\overline{\Phi}_{1}^{k_{\rm c}}(M+1,2)=c^{2}\sum_{j=1}^{k-M-2}(k-1-j)^{-2}(j+M)^{-2}=c^{2}\sum_{j=M+1}^{k-2}(k+M-1-j)^{-2}j^{-2}$
$\leq$ $2c^{2}(k+M-1)^{-2}\sum_{j=M+1}^{k-2}\big((k+M-1-j)^{-2}+j^{-2}\big)\leq 4c_{K,1}^{2}c^{2}M^{-1}(k+M)^{-2}.$
@@
$(j^{\prime}-j)^{-3}j^{-2}=3(j^{\prime})^{-4}\big((j^{\prime}-j)^{-1}+j^{-1}\big)+(j^{\prime})^{-3}\big(2(j^{\prime}-j)^{-2}+j^{-2}\big)+(j^{\prime})^{-2}(j^{\prime}-j)^{-3},$
@@
$\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)=c^{2}\sum_{j=1}^{k-M-2}(k-1-j)^{-3}(j+M)^{-2}=c^{2}\sum_{j=M+1}^{k-2}(k+M-1-j)^{-3}j^{-2}$
$\leq$ $c^{2}\big(k+M-1\big)^{-2}\bigg[6\big(k+M-1\big)^{-2}\sum_{j=M+1}^{k-2}j^{-1}+3\big(k+M-1\big)^{-1}\sum_{j=M+1}^{k-2}j^{-2}+\sum_{j=M+1}^{k-2}j^{-3}\bigg]$
$\leq$ $c^{2}\big(k+M-1\big)^{-2}\Big[6\big((K+1)M\big)^{-2}\ln(KM+t-1)+3\big((K+1)M\big)^{-1}M^{-1}+\tfrac{1}{2}M^{-2}\Big]$
$\leq$ $c_{K,1}^{2}c^{2}(k+M)^{-2}M^{-2}\Big[6(K+1)^{-2}\big(\ln(K+1)+\ln M\big)+\tfrac{3}{K+1}+\tfrac{1}{2}\Big].$
@@
implies the bound on $\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)$. For the estimate (A.10), the splitting $(j^{\prime}-j)^{-1}j^{-2}=(j^{\prime})^{-1}j^{-2}+(j^{\prime})^{-2}\big((j^{\prime}-j)^{-1}+j^{-1}\big)$ implies
$\overline{\Phi}_{1}^{k-1}(0,1)=c^{2}\sum_{j=1}^{k-1}(k-1-j)^{-1}(j+M)^{-2}=c^{2}\sum_{j=M+1}^{k+M-1}(k+M-1-j)^{-1}j^{-2}$
$\leq$ $c^{2}(k+M-1)^{-1}\bigg[\sum_{j=M+1}^{k+M-1}j^{-2}+(k+M-1)^{-1}\sum_{j=M+1}^{k+M-1}\big((k+M-1-j)^{-1}+j^{-1}\big)\bigg]$
$\leq$ $c^{2}(k+M-1)^{-1}\big[M^{-1}+(k+M-1)^{-1}\big(2+2\ln(k+M-1)\big].$

### v2

@@ 5 Concluding remarks
In this work, we have investigated stochastic variance reduced gradient (SVRG) and a regularized variant (rSVRG) for solving linear inverse problems in Hilbert spaces. We have established the regularizing property of both SVRG and rSVRG. Under the source condition, we have derived convergence rates in expectation and in the uniform sense for (r)SVRG. These results indicate the optimality of SVRG for nonsmooth solutions and the built-in regularization mechanism and optimality of rSVRG. The numerical results for three linear inverse problems with varying degree of ill-posedness show the advantages of rSVRG over SGD, standard SVRG and Landweber method. Note that both SVRG and rSVRG depend on the knowledge of the noise level, and the a priori choice of the relevant algorithmic parameters also require a knowledge of the unknown regularity index $\nu$. Thus there is a need to develop data-driven or adaptive rules for choosing these algorithmic parameters. Moreover, in practice, the noise level may be unknown, and certain heuristic techniques are required for their efficient implementation, e.g., as the a priori stopping rule or constructing the approximate operator $A$. We leave these interesting questions to future works.
Acknowledgments
The authors are grateful to the editor and two anonymous referees for their constructive comments which have improved the quality of the paper.
@@ Acknowledgments
The first inequality can be found in [17, Lemma 3.4]. To show the second inequality, let $\mathrm{Sp}(P)$ be the spectrum of $P$. Then there holds
@@
$\|B^{\frac{1}{2}}(I-P^{t})P^{k}\|$ $\leq\|B\|^{\frac{1}{2}-\epsilon}\|B^{\epsilon}P^{\frac{k+t}{2}}\|\|(I-P^{t})P^{\frac{k-t}{2}}\|$
$\leq 2^{1+\epsilon}\epsilon^{\epsilon}c_{0}^{-\epsilon}\|B\|^{\frac{1}{2}-\epsilon}t(k+t)^{-(1+\epsilon)}.$
@@
$\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=\mathbb{E}[\|R(B-A_{i_{j}}^{*}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
$=$ $\mathbb{E}[\|RA_{i_{j}}^{*}A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]-\|RB\Delta_{j}^{\delta}\|^{2}\leq L\|R\|^{2}\mathbb{E}[\|A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
$=$ $L\|R\|^{2}n^{-1}\sum_{i=1}^{n}\|A_{i}\Delta_{j}^{\delta}\|^{2}=n^{-1}L\|R\|^{2}\|A\Delta_{j}^{\delta}\|^{2}.$
@@
$\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]=\mathbb{E}[\|RA^{*}(n^{-1}A-b_{i_{j}}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
$\leq$ $\mathbb{E}[\|RA^{*}\|^{2}\|(n^{-1}A-b_{i_{j}}A_{i_{j}})\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
$=$ $n\|RB^{\frac{1}{2}}\|^{2}\big(\mathbb{E}[\|b_{i_{j}}A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]-\|n^{-1}A\Delta_{j}^{\delta}\|^{2}\big)$
$\leq$ $n\|RB^{\frac{1}{2}}\|^{2}\mathbb{E}[\|A_{i_{j}}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]\leq\|RB^{\frac{1}{2}}\|^{2}\|A\Delta_{j}^{\delta}\|^{2}.$
These estimates and the inequality
$\|RN_{j}\Delta_{j}^{\delta}\|^{2}\leq n\mathbb{E}[\|RN_{j}\Delta_{j}^{\delta}\|^{2}|\mathcal{F}_{j}]$
complete the proof of the lemma. ∎
The proof of Theorem 3.1 is lengthy and very technical, and requires several technical lemmas with sharp constants. The first lemma provides useful bounds on the bias and variance components of the weighted successive error $A\Delta_{k}^{\delta}$ in terms of the iteration index $k$.
@@
Let $k=KM+t$ with $K\geq 0$ and $1\leq t\leq M-1$. Similar to the proof of Lemma 3.2, for the bias term $\|\mathbb{E}[A\Delta_{KM+t}^{\delta}]\|$, by the definitions of $\zeta$, $B$ and $\tilde{\phi}^{t-1}$, the identity
@@
Next let
$S_{t,j}=c_{0}\|A(P^{t}-I)P^{KM-1-j}N_{j}\Delta_{j}^{\delta}\|\quad\mbox{and}\quad T_{t,j}=c_{0}\|AP^{KM+t-1-j}N_{j}\Delta_{j}^{\delta}\|.$
Then for the variance, when $K\geq 1$, by Lemma 3.1 and the identity $\mathbb{E}[\langle R_{1}N_{i}\Delta_{i}^{\delta},R_{2}N_{j}\Delta_{j}^{\delta}\rangle|\mathcal{F}_{j}]=0$ for any $j>i$ and deterministic bounded linear operators $R_{1}$ and $R_{2}$, we have
@@
with the three terms ${\rm I}_{1}$, ${\rm I}_{2}$ and ${\rm I}_{3}$ given by
${\rm I}_{1}=\sum_{j=1}^{(K-1)M+t-2}\mathbb{E}[S_{t,j}^{2}],\quad{\rm I}_{2}=\sum_{j=(K-1)M+t-1}^{KM-1}\mathbb{E}[S_{t,j}^{2}]\quad\mbox{and}\quad{\rm I}_{3}=\sum_{j=KM+1}^{KM+t-1}\mathbb{E}[T_{t,j}^{2}].$
@@
$\mathbb{E}[S_{t,j}^{2}]$ $\leq c_{0}^{2}L\|B^{\frac{1}{2}}(P^{t}-I)P^{KM-1-j}\|^{2}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]$
$\leq c_{0}^{2}L\|B\|\|(P^{t}-I)P^{KM-1-j}\|^{2}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
@@
${\rm I}_{2}\leq$ $c_{0}^{2}L\|B\|\sum_{j=(K-1)M+t-1}^{KM-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
${\rm I}_{3}\leq$ $c_{0}^{2}L\|B\|\sum_{j=KM+1}^{KM+t-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}].$
Meanwhile, by the commutativity of $B,P$ and Lemma A.1 with $\epsilon=\frac{1}{2}$, we can alternatively get
@@
${\rm I}_{2}\leq$ $\frac{c_{0}L}{2}\left(2c_{0}\|B\|\mathbb{E}[\|A\Delta_{KM-1}^{\delta}\|^{2}]+\sum_{j=(K-1)M+t-1}^{KM-2}(KM-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\right),$
${\rm I}_{3}\leq$ $\frac{c_{0}L}{2}\left(2c_{0}\|B\|\mathbb{E}[\|A\Delta_{KM+t-1}^{\delta}\|^{2}]+\sum_{j=KM+1}^{KM+t-2}(KM+t-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\right).$
@@
$\mathbb{E}[\|A\Delta_{t}^{\delta}-\mathbb{E}[A\Delta_{t}^{\delta}]\|^{2}]=\sum_{j=1}^{t-1}\mathbb{E}[T_{t,j}^{2}]$
$\leq$ $\left\{\begin{aligned} c_{0}^{2}L\|B\|\sum_{j=1}^{t-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],\\ \frac{c_{0}L}{2}\left(2c_{0}\|B\|\mathbb{E}[\|A\Delta_{t-1}^{\delta}\|^{2}]+\sum_{j=1}^{t-2}(t-1-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\right).\end{aligned}\right$.
Then combining the preceding estimates with the identity $\|B\|=n^{-1}\|A\|^{2}$ and the inequality $t\leq M-1$ gives the estimate (A.2). Finally, when $K\geq 1$, by Lemma 3.1 and the triangle inequality, we derive
$\|A\Delta_{KM+t}^{\delta}-\mathbb{E}[A\Delta_{KM+t}^{\delta}]\|\leq$ $\sum_{j=1}^{KM-1}S_{t,j}+\sum_{j=KM+1}^{KM+t-1}T_{t,j}.$
@@
Combining these estimates with the inequality $t\leq M-1$ gives the estimate (A.3). ∎
The next lemma gives several basic estimates on the following summations:
$\overline{\Phi}_{j_{1}}^{j_{2}}(0,1)$ $=2c_{0}\|B\|\mathbb{E}[\|A\Delta_{j_{2}}^{\delta}\|^{2}]+\sum_{j=j_{1}}^{j_{2}-1}(j_{2}-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
$\Phi_{j_{1}}^{j_{2}}(0,\tfrac{1}{2})$ $=\sqrt{2c_{0}\|B\|}\|A\Delta_{j_{2}}^{\delta}\|+\sum_{j=j_{1}}^{j_{2}-1}(j_{2}-j)^{-\frac{1}{2}}\|A\Delta_{j}^{\delta}\|,$
$\overline{\Phi}_{j_{1}}^{j_{2}}(i,r)$ $=\sum_{j=j_{1}}^{j_{2}}(j_{2}+i-j)^{-r}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}],$
$\Phi_{j_{1}}^{j_{2}}(i,r)$ $=\sum_{j=j_{1}}^{j_{2}}(j_{2}+i-j)^{-r}\|A\Delta_{j}^{\delta}\|,$
for any r,i>0r,\;i>0 or $r=0$. The proof of these results is elementary but lengthy and tedious in order to get precise constants.
@@
with $K=[\frac{k}{M}]$, and the constants $c_{K,1}$, $c_{K,2}$ and $c_{K,\epsilon}$ defined respectively by
$c_{K,1}$ $=1+\frac{2}{K}\leq 3,$
$c_{K,2}$ $=2+\frac{3}{K+1}+\frac{6\ln M}{(K+1)^{2}}\leq\frac{7}{2}+\frac{3}{2}\ln M,$
$c_{K,\epsilon}$ $=\frac{e^{-1}+1}{\epsilon}+\frac{2^{\epsilon}\ln M}{(K+1)^{\epsilon}}.$
Note that there hold $\lim_{K\to\infty}c_{K,1}=1$, $\lim_{K\to\infty}c_{K,2}=2$, and $\lim_{K\to\infty}c_{K,\epsilon}=(e^{-1}+1)\epsilon^{-1}$.
@@
Let $k=KM+t$ with $K\geq 1$ and $t=1,\cdots,M-1$. In the proof, we use the shorthand notation $\overline{k}=k-1$ and $\overline{k_{M}}=k_{M}-1$. Then there holds the elementary inequality:
$\overline{k}^{-1}\leq c_{K,1}(k+M)^{-1}.$ (A.12)
@@
$\overline{\Phi}_{k_{\rm c}+1}^{\overline{k}}(0,0)$ $=\sum_{j=k_{\rm c}+1}^{\overline{k}}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]\leq c^{2}M\overline{k}^{-2}\leq c_{K,1}^{2}c^{2}M(k+M)^{-2},$
$\Phi_{k_{\rm c}+1}^{\overline{k}}(0,0)$ $=\sum_{j=k_{\rm c}+1}^{\overline{k}}\|A\Delta_{j}^{\delta}\|\leq cM\overline{k}^{-1}\leq c_{K,1}cM(k+M)^{-1}.$
@@
$\overline{\Phi}_{k_{\rm c}+1}^{\overline{k_{M}}}(0,1)+\overline{\Phi}_{k_{M}+1}^{\overline{k}}(0,1)$
$=$ $\sum_{j=k_{\rm c}+1}^{k_{M}-2}(\overline{k_{M}}-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]+\sum_{j=k_{M}+1}^{k-2}(\overline{k}-j)^{-1}\mathbb{E}[\|A\Delta_{j}^{\delta}\|^{2}]$
$+2c_{0}\|B\|\big(\mathbb{E}[\|A\Delta_{\overline{k_{M}}}^{\delta}\|^{2}]+\mathbb{E}[\|A\Delta_{\overline{k}}^{\delta}\|^{2}]\big)$
$=$ $\sum_{j=1}^{M-t}j^{-1}\mathbb{E}[\|A\Delta_{\overline{k_{M}}-j}^{\delta}\|^{2}]+\sum_{j=1}^{t-2}j^{-1}\mathbb{E}[\|A\Delta_{\overline{k}-j}^{\delta}\|^{2}]$
$+2c_{0}\|B\|\big(\mathbb{E}[\|A\Delta_{\overline{k_{M}}}^{\delta}\|^{2}]+\mathbb{E}[\|A\Delta_{\overline{k}}^{\delta}\|^{2}]\big)$
$\leq$ $c^{2}\bigg(\sum_{j=1}^{M-t}j^{-1}(\overline{k_{M}}-j+M)^{-2}+\sum_{j=1}^{t-2}j^{-1}(\overline{k}-j+M)^{-2}$
$\quad+2c_{0}\|B\|\big((\overline{k_{M}}+M)^{-2}+(\overline{k}+M)^{-2}\big)\bigg)$
$\leq$ $c^{2}\overline{k}^{-2}{\rm I},$
with
${\rm I}=4c_{0}\|B\|+\sum_{j=1}^{M-t}j^{-1}+\sum_{j=1}^{t-2}j^{-1}.$
Note that the factor $\rm I$} can be bounded by
${\rm I}$ $\leq 4c_{0}\|B\|+2+\ln(M-t)+\ln t\leq 4c_{0}\|B\|+2+2\ln\tfrac{M}{2}$
$=4c_{0}\|B\|+2-2\ln 2+2\ln M\leq 1+2\ln M+4c_{0}\|B\|.$
@@
$\overline{\Phi}_{k_{\rm c}+1}^{\overline{k_{M}}}(0,1)+\overline{\Phi}_{k_{M}+1}^{\overline{k}}(0,1)\leq c_{K,1}^{2}c^{2}(1+2\ln M+4c_{0}\|B\|)(k+M)^{-2}.$
@@
$\overline{\Phi}_{1}^{k_{\rm c}}(M+1,2)\leq c^{2}\sum_{j=1}^{k-M-2}(\overline{k}-j)^{-2}(j+M)^{-2}=c^{2}\sum_{j=M+1}^{k-2}(\overline{k}+M-j)^{-2}j^{-2}$
$\leq$ $2c^{2}(\overline{k}+M)^{-2}\sum_{j=M+1}^{k-2}\big((\overline{k}+M-j)^{-2}+j^{-2}\big)\leq 4c_{K,1}^{2}c^{2}M^{-1}(k+M)^{-2}.$
@@
$(j^{\prime}-j)^{-3}j^{-2}=$ $3(j^{\prime})^{-4}\big((j^{\prime}-j)^{-1}+j^{-1}\big)$
$+(j^{\prime})^{-3}\big(2(j^{\prime}-j)^{-2}+j^{-2}\big)+(j^{\prime})^{-2}(j^{\prime}-j)^{-3},$
@@
$\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)\leq c^{2}\sum_{j=1}^{k-M-2}(\overline{k}-j)^{-3}(j+M)^{-2}=c^{2}\sum_{j=M+1}^{k-2}(\overline{k}+M-j)^{-3}j^{-2}$
$=$ $c^{2}\big(\overline{k}+M\big)^{-2}\bigg[6\big(\overline{k}+M\big)^{-2}\sum_{j=M+1}^{k-2}j^{-1}+3\big(\overline{k}+M\big)^{-1}\sum_{j=M+1}^{k-2}j^{-2}+\sum_{j=M+1}^{k-2}j^{-3}\bigg]$
$\leq$ $c^{2}\big(\overline{k}+M\big)^{-2}\Big[6\big((K+1)M\big)^{-2}\ln(KM+t-1)+3\big((K+1)M\big)^{-1}M^{-1}+\tfrac{1}{2}M^{-2}\Big]$
$\leq$ $\frac{c_{K,1}^{2}c^{2}}{(k+M)^{2}M^{2}}\Big[6(K+1)^{-2}\big(\ln(K+1)+\ln M\big)+\tfrac{3}{K+1}+\tfrac{1}{2}\Big].$
@@
implies the desired bound on $\overline{\Phi}_{1}^{k_{\rm c}}(M+1,3)$. For the estimate (A.10), the splitting
$(j^{\prime}-j)^{-1}j^{-2}=(j^{\prime})^{-1}j^{-2}+(j^{\prime})^{-2}\big((j^{\prime}-j)^{-1}+j^{-1}\big)$
implies
$\overline{\Phi}_{1}^{\overline{k}}(0,1)\leq c^{2}\Big(2c_{0}\|B\|(\overline{k}+M)^{-2}+\sum_{j=1}^{k-2}(\overline{k}-j)^{-1}(j+M)^{-2}\Big)$
$=$ $c^{2}\Big(2c_{0}\|B\|(\overline{k}+M)^{-2}+\sum_{j=M+1}^{k+M-2}(\overline{k}+M-j)^{-1}j^{-2}\Big)$
$=$ $c^{2}(\overline{k}+M)^{-1}\bigg[\sum_{j=M+1}^{k+M-2}j^{-2}+(\overline{k}+M)^{-1}\Big(2c_{0}\|B\|+\sum_{j=M+1}^{k+M-2}\big((\overline{k}+M-j)^{-1}+j^{-1}\big)\Big)\bigg]$
$\leq$ $c^{2}(\overline{k}+M)^{-1}\big[M^{-1}+(\overline{k}+M)^{-1}\big(1+2\ln(\overline{k}+M)+2c_{0}\|B\|\big)\big].$